what is the slope of a line containing the points (-2,-4) and (4,5)
Answer by lenny460(1073)

What is the slope of a line containing the points:
(-2,-4) and
(4,5)
Slope (m) = (y2 - y1) / (x2 - x1)
= 5 - (-4) / 4 - (-2)
= 5 + 4 = 9
= 4 + 2 = 6
The Slope:
9/6 = 3/2 = 1 1/2
Second Method:
Slope = rise/run
When you draw the graph:
rise = 9
run = 6
Slope:
9/6 = 3/2 = 1 1/2